May you please help me with this problem:
5e+4f=9
4e+5f=9
The instructions are to use linear combinations to solve the linear system. Then check your solution. Thank you for your time.

Answer by Alan3354(69443)

Do it like this one.
------------
3j+5k=19
j-2k=-1
------------
Multiply the 2nd eqn by 3 to make the coefficients equal
3j+5k=19
3j-6k=-3
---------- Subtract
-k = 22
k = 22
-------
j - 2k = -1
j -44 = -1
j = 43